Qd= 25000-2p








Qs= 10000-1p








Calculate the market equlibrium level of price and quantity for housing unit?


Calculate price elasticity of demand using point elasticity method when industry is in equlibrium?


1
Expert's answer
2021-05-05T13:31:42-0400

An equilibrium is that state of balance

"Qd=25000-2P"

"Qs=10000+1P"

Equilibrium is found out by equating Demand = Supply.

So, 


"25000-2P=10000+1P"

"25000+10000=1P+2P"

"35000=3P"

"P=\frac{35000}{3}"

"P=11,666.67"


"Qd=25000-2(11,666.67)"

"Qd=25000-23333.33"


Equilibrium Price = 11666.67

Equilibrium Quantity = 1666.67


Elasticity is the change in quantity due to change in price. 

"ed=\frac{\delta Q}{\delta P}\times \frac{P}{Q}"


"ed=\frac{\delta (25000-2P)}{\delta P}\times \frac{11666.67}{1666.67}"


"ed=-2\times 6.99988"

"ed=-13.99"

Price Elasticity of Demand using point elasticity method is 13.99. That is demand is highly elastic.